OpenStudy (anonymous):

Continuous random variable X has density given by f(x)=c⋅(x+1)⋅(x−3) if x∈[−1,3] and 0 otherwise. 1) Find DF(2;X) . 2) Find E[X] .

OpenStudy (anonymous):

\[f(x)=\begin{cases}c(x+1)(x-3)&\text{for }x\in[-1,3]\\[1ex]0&\text{otherwise}\end{cases}\]In order for \(f(x)\) to be a valid pdf you must have \[\int_{-\infty}^\infty f(x)\,\mathrm{d}x=c\int_{-1}^3(x+1)(x-3)\,\mathrm{d}x=1\]Finding \(c\) isn't too hard. I'm not familiar with the notation \(\mathrm{DF}(2;X)\). I assume \(\mathrm{DF}\) stands for "distribution function", ie. cdf, but I'm not sure what the \(2;X\) is supposed to mean. Maybe the value of the cdf at \(X=2\)? The expected value is simple enough once you determine the value of \(c\): \[\mathbb{E}(X)=\int_{-\infty}^\infty xf(x)\,\mathrm{d}x=c\int_{-1}^3x(x+1)(x-3)\,\mathrm{d}x=\cdots\]